OpenStudy (anonymous):

I need help with a curve dissection wkst. There are 12 problems. Some of them I need help with and others I just want to make sure I did them right. (AP Calculus)

OpenStudy (anonymous):

We are instructed to find the critical points by using the first derrivative. Then find the points of infliction using the second derrivative. Then to draw the curve on a graph

OpenStudy (anonymous):

the first problem is \[y=3x ^{2/3}-2x\] the derrivative is (i think ) \[y \prime = 2x ^{1/3}-2\] which means that the critical point is x=0? the second derrivative is (once again, i am not positive) \[y \prime = \frac{ 2 }{ 3 }x ^{2/3}\] makng the point of infliction x=0?

OpenStudy (anonymous):

Remember that \[ \frac{ d(x^n) }{ dx } = nx^{n-1}\] you only decremented n by \[\frac{1}{3}\]

OpenStudy (anonymous):

so the derrivative would be \[y \prime = 2x ^{-1/3}-2\]

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

how woudi find he critical pt for that?

OpenStudy (anonymous):

i know the derrivative is also equal to\[y \prime = 2(x ^{-1/3}-1)\] if that helps any

OpenStudy (anonymous):

set y prime to 0 and solve as you would normally

OpenStudy (anonymous):

ok. so i get. . .\[0= 2x ^{-1/3}-2\] add 2 to other side \[2x ^{-1/3}=2\] divide by 2 \[x ^{-1/3}=1\] then i get stuck

OpenStudy (anonymous):

well\[x^{-1/3} = \frac{1}{x^{1/3}} = 1\] right? try cubing both sides!

OpenStudy (anonymous):

did this help?